### v1

@@ 4.4 Ablation Study
Tab. 4, Fig. 6 shows the results of our comparative experiments on the efficiency of GSAM. As a comparison method, the MACs values of SAM, LoRA, ConvLoRA, AdaptFormer, and GSAM are compared when the size of the random cropping is changed. Since the number of input image sizes is fixed in conventional SAM fine-tuning methods, the computational cost becomes huge, and it can be seen that the computational cost is more than 300G MACs for all the methods except SAMUS. On the other hand, with GSAM, the computational cost decreases exponentially as the size of the random crop is reduced. In particular, when the input image size is $128\times 128$, the segmentation accuracy outperforms all conventional methods, despite a computational cost of around half that of SAMUS. The SM-Adaptformer included in GSAM has a relatively complex structure to acquire multi-scale features and improve segmentation accuracy, but by using PEG and CNN to support variable-size input, the input image size can be reduced, which reduces the computational cost has been significantly reduced. Based on these results, we expect GSAM to be widely used as an efficient approach for fine-tuning SAM in the future, as it significantly reduces the computational cost and allows for highly accurate segmentation.

### v2

@@ 4.4 Ablation Study
Tab.˜4, Fig.˜6 shows the results of our comparative experiments on the efficiency of GSAM. As a comparison method, the MACs values of SAM, LoRA, ConvLoRA, AdaptFormer, and GSAM are compared when the size of the random cropping is changed. Since the number of input image sizes is fixed in conventional SAM fine-tuning methods, the computational cost becomes huge, and it can be seen that the computational cost is more than 300G MACs for all the methods except SAMUS. On the other hand, with GSAM, the computational cost decreases exponentially as the size of the random crop is reduced. In particular, when the input image size is $128\times 128$, the segmentation accuracy outperforms all conventional methods, despite a computational cost of around half that of SAMUS. The SM-Adaptformer included in GSAM has a relatively complex structure to acquire multi-scale features and improve segmentation accuracy, but by using PEG and CNN to support variable-size input, the input image size can be reduced, which reduces the computational cost has been significantly reduced. Based on these results, we expect GSAM to be widely used as an efficient approach for fine-tuning SAM in the future, as it significantly reduces the computational cost and allows for highly accurate segmentation.